On December 31, 2003 AG Inc. had a workingcapital ratio (current ratio) of 2, and reported the following accounts; Cash $13,000 Accounts Payable $2,000 Accounts Receivable $4,000 Wages Payable $5,500 Allowance for Uncollectible Accounts ? Utilities Payable $500 Inventory $6,600 Interest Payable $4,000 Prepaid Insurance $600 Bonds Payable $13,000 Equipment (net) $24,000 Retained Earnings $3,600 Therefore, the balancein the Allowance for Uncollectible Accountswas:

$200

$300

$600

None of the other alternatives are correct

$0
Answer #1
Therefore, the balancein the Allowance for Uncollectible Accountswas $200
Dear Student
Thank you for using Chegg
Please find below the answer
Statementshowing Computations
Paticulars Amount
Current Assets:
Cash            13,000.00
Accounts Receivable 4000 - x
Inventory              6,600.00
Prepaid Insurance                  600.00
Current Assets 24,200 - x
Current Liabilities:
Accounts Payable              2,000.00
Wages Payable              5,500.00
Utilities Payable                  500.00
Interest Payable              4,000.00
Current Liabilities            12,000.00
Current Ratio = CA/CL
2 = 24,200 -x /12,000
24,000 = 24,200 -x
x= 200